…..........................................How do you solve the sum of  tan 27 X tan 63
 
 

Grade:8

1 Answers

Aditya Gupta
5 years ago
so we know that tany= cot(90-y)=1/tan(90-y)
if we let y=27, then
tan27=1/tan(90-27)=1/tan63
or tan27*tan63= 1.